# LLM Collaboration with Multi-Agent Reinforcement Learning

### v2

@@ Appendix B Cooperation Schemes
B.1 Fallback
The most commonly observed cooperation scheme is the main agent providing a fallback for the auxiliary function. Although prompted to use the auxiliary output, the main agent may still doubt its completeness and reimplement the solution as a backup in case of potential errors.
@@ B.2 Decorator
4 else:
5 return aux1(text[1:-1])
6
7def aux2(text):
@@
10 else:
11 return aux2(text[1:-1])
12
13def is_palindrome(text):
@@
18 return text[0] == text[-1] and aux1(text)
19
@@
21 return text[0] == text[-1] and aux2(text)
B.3 Coordinator
@@ B.4 Strategy Filter
Appendix C Broader Impacts
@@
Prompt-based coordination is often brittle Estornell and Liu [2024], as agents may fail to follow instructions they were not explicitly trained to interpret. Our method builds on a solid theoretical foundation in cooperative MARL, explicitly optimizing agents for joint optimality. Our work also opens opportunities to enhance existing test-time multi-agent interaction methods by integrating MARL techniques Du et al. [2023], Lifshitz et al. [2025], Wu et al. [2023a], particularly in settings that involve task decomposition and iterative feedback integration.
@@
This work also explores a new perspective on accelerating LLM inference through cooperative MARL. While mainstream acceleration techniques (e.g., knowledge distillation, pruning, and quantization) improve efficiency at the cost of information loss Wang et al. [2024], Zhao et al. [2024], our approach suggests decentralized coordination among specialized agents, thereby alleviating the burden of long-context memory and joint decision-making on a single model. Each agent can focus on a specific subtask, enabling more modular and robust reasoning.
Appendix D Limitations and Future Works
@@
Nevertheless, this study is subject to several limitations. First, we focus on homogeneous agents for simplicity, assuming they perform similar tasks despite being assigned different roles, e.g., both the auxiliary agent and main agent are generating Python functions. Future research could explore LLM collaboration among heterogeneous agents with diverse capabilities and functionalities.
@@
Due to computational constraints, we train LLMs with MAGRPO on limited datasets using relatively small-scale language models. When LLM-based coding agents are deployed in larger-scale projects involving multiple files and modules, more diverse and complex cooperation schemes are likely to emerge, which would further demonstrate the potential of decentralized coordination in MAS.
@@
The simplicity of our reward model inevitably leads to narrow reward signals and potential reward hacking. As suggested by many research studies and industrial practice Uesato et al. [2022], Wu et al. [2023b], designing more expressive and fine-grained reward models (e.g., multi-aspect rewards, process-supervised rewards) is essential for better aligning agent cooperation with human preferences.

### v3

@@ Appendix B Additional Results
Fallback
A commonly observed cooperation scheme is the main agent providing a fallback for the auxiliary function. Although prompted to use the auxiliary output, the main agent may still doubt its completeness and reimplement the solution as a backup in case of potential errors.
Decorator
@@
4 return text[0] == text[-1] and aux1(text[1:-1])
5
6def aux2(text):
@@
9 return text[0] == text[-1] and aux2(text[1:-1])
10
11def is_palindrome(text):
@@
16 return aux1(text)
17
@@
19 return aux2(text)
@@
Learning Modes
Figure 5 shows the reward curves of total returns in a 2-turn MAGRPO training with 2 different learning modes.
(a) Self-Improvement
(b) Guided-Learning
Figure 5: Learning modes in 2-turn MAGRPO training.
Figure 5(a) demonstrates a self-improvement mode, where agents interact with the tasks themselves and do not have access to the external feedback. At each turn $t$, the prompts are just prompts and responses from the previous turns and a revision instruction. Although both curves improve as agents gradually develop cooperative behaviors, the performance of the second turn is still consistently similar to the first turn, suggesting that learning is primarily driven by direct task interaction.
Figure 5(b) illustrates the reward curves of guided-learning mode, where LLMs leverage external feedback to improve performance, e.g., feedback from expert agents, diagnosis from static checkers like AST, and sandbox tests. When employing Claude-Sonnet-4 as an external model to provide more concrete suggestions or feeding error messages from sandbox tests, the performance of the second turn (red) exceeds first turn (blue), and both outperform those in the self-improvement. This indicates that appropriate guidance helps agents to refine the response in an efficient way. Due to the computational constraints, most models used in our setup have around 3B parameters and may struggle to interpret more abstract suggestions. We hypothesize that larger models with higher reasoning capabilities could benefit from more implicit guidance.
Appendix C Experimental Configurations
This section outlines the hyperparameter settings and reward specifications used in our experiments.
Hyperparameters
For writing collaboration, we set the temperature to $0.8$ and apply nucleus sampling with a threshold of $0.95$ to encourage diverse generation. Policy deviation is regularized using a beta value of $0.02$. The policy is optimized using a learning rate of $5\times 10^{-6}$, and training is conducted for 1,500 steps.
For coding collaboration, the single-turn MAGRPO training uses a temperature of $0.7$ and nucleus sampling with a threshold of $0.9$. The learning rate is set to $1\times 10^{-6}$, with 1,500 training steps. In the multi-turn setting, the discount factor is set to $1.0$, and the learning rate is $5\times 10^{-6}$, with 2,200 training steps.
Reward Specifications
Rewards are computed as a weighted sum of multiple metric-based components, following a hierarchical reward modeling scheme to prioritize fundamental objectives.
TLDR Summarization
[…]
  • Style Consistency: The style consistency is measured through Jaccard similarity of vocabulary between the completions (excluding stopwords). The Jaccard similarity scores are capped at $0.03$ and normalized as rewards. We use a cap here to balance the needs of maintaining lexical consistency and vocabulary expansion in elaborated summarization.
  • Logical Coherence: The logical coherence is evaluated through the presence and diversity of transition words in the completions. We check transition words across 12 categories, e.g., examples, explanation, contrast, etc. Additional rewards are given for using transition words in more categories, where $r=\min(0.6\log(n+1),1)$, and $n$ is the number of transition categories.
arXiv Expansion
[…]
  • Style Consistency: Style consistency is quantified using Jaccard similarity between the 2 completions. The similarity score is capped at 0.23 and normalized as rewards.
  • Logical Coherence: Logical coherence is assessed based on the presence of transition words across 12 categories. Additional rewards are given for using transition words in more categories, where $r=\min(0.4\log(n+1),1)$, and $n$ is the number of transition categories.
Coding Collaboration
  • Structural Integrity: This metric verifies the correct implementation of both the auxiliary and main functions. To receive the base reward, the corresponding functions in the agents’ completions must be properly defined and include return statements. Failure to define the main function results in evaluation termination.
  • Syntactical Correctness: This metric assesses the syntactic validity of the concatenated code, which includes the libraries provided in the dataset, the auxiliary function defined by the helper agent, and the function defined by the main agent. Syntactic correctness is verified via static analysis, i.e., Abstract Syntax Tree (AST). The presence of syntax errors leads to the termination of the evaluation to avoid runtime failures.
  • Test Pass Rate: This metric measures the percentage of unit tests passed during execution, with each test subject to a 10-second timeout. Rewards are assigned proportionally based on the number of successful assertions. If no tests pass, the evaluation is terminated.
  • Cooperation Quality: A base bonus is applied if the main function calls the auxiliary. Additional rewards are given when the main function implements substantive logic beyond simply wrapping the auxiliary.
Appendix D Prompt Design
Writing Collaboration
TLDR
In the TLDR summarization, the prompt field of the dataset is processed for each agent by using the following instructions.
Summary Agent
Create a concise summary response to this post.
Query: {prompt}
Instructions: Provide a brief and focused summary in a few sentences
Elaboration Agent
Create a detailed summary response to this post.
Query: {prompt}
Instructions: You should use transition words to improve flow
arXiv
In the arXiv paragraph expansion, we use the abstract field of the dataset and process it as follows.
Background Agent
Based on the following scientific abstract, expand the content for the introduction section.
Abstract: {abstract}
Instructions:
- There is another agent that will provide the method and implications
- You just need to focus on the background and motivation
- Avoid repeating methodology and implications content
Method Agent
Based on the following scientific abstract, expand the content for the introduction section.
Abstract: {abstract}
Instructions:
- There is another agent that will provide the background and motivation
- You just need to focus on the method and implications
- Avoid repeating background and motivation content
Coding Collaboration
For HE and CHE, we extract the entry_point, params from the prompt field and instruct agents as follows.
Auxiliary Agent
Create a helper function for this coding problem.
Problem: {prompt}
Instructions:
- Output ONLY the function code, no explanations or examples
- Do NOT include markdown code blocks (‘‘‘python)
- Do NOT include any text before or after the function
- Do NOT include test cases or example usage
- Create a helper function named aux that can assist the main function
- The function should return useful data for solving the problem
Your output should follow this format:
def aux(...):
# your code here
return result
Main Agent
Solve this coding problem by implementing the required function.
Problem: {prompt}
You have access to a helper function: aux(...)
Instructions:
- Output ONLY the function code, no explanations or examples
- Do NOT include markdown code blocks (‘‘‘python)
- Do NOT include any text before or after the function
- Do NOT include test cases or example usage
- Do NOT redefine the aux() function
- Implement ONLY the ’{entry_point}’ function as specified
- You can call aux() to assign a value to a variable within your function if helpful
Your output should follow this format:
[…]
To improve the generated code, these prompts are used to construct second-turn observations for the MAS with suggestions from Claude-Sonnet-4.
External Agent
You are an advisor helping 2 agents (an auxiliary agent and a main agent) solve the following problem. The auxiliary agent provides a helper function (aux), while the main agent defines the task-specific logic.
Problem: {prompt}
Example tests: {test}
Show your feedback and edits for the following code: {combined_code}
Instructions:
- If you identify a missing element, such as an undefined aux or missing entry point (main function), you should write one for it.
- If both are not missing, point out and make changes to any critical syntax or logic errors that would prevent the code from passing the given unit tests.
- You should focus only on clear errors on the given unit tests, be conservative and lenient, ignoring issues like redundancy, inefficiency, lack of edge case handling, or type annotations.
- Return "Perfect! No edits needed!" if the logic is sound.
Your response MUST contain the JSON format specified below. Always include both aux and main fields, even if no edits are needed.
{ "aux": {{...}}, "main": {{...}}}
Appendix E Baseline Methods
We adopt a single-agent method and 3 representative multi-agent conversation methods as baselines.
Figure 6: Single-agent code generation.
Figure 6 illustrates the code generation process using a single LLM agent. In this setting, the user gives a coding question along with specific instructions. The agent responds by generating a Python function snippet to solve it.
Figure 7: Coding collaboration via naive concatenation.
Naive concatenation represents the simplest form of cooperation, where 2 agents generate code synchronously, as illustrated in Figure 7. The first agent is provided with the coding question and informed of its role as a helper. The second agent is given the same question, along with its role as the main generator and the fact that an auxiliary agent exists. However, the main agent lacks information about the specific functionality of the auxiliary agent. Their outputs are directly concatenated to form the response. This method is intended to improve inference efficiency by enabling a simple division of the problem into separate parts.
Figure 8: Coding collaboration via sequential pipeline.
Figure 8 presents the form of pipeline cooperation, where agents respond in sequence. The first agent is given the coding question along with the role of a helper. Its response is then passed to the main agent as a reference. The main agent generates its answer by incorporating the helper’s response. This method enables one-way communication, allowing the main agent to respond by coordinating with the helper. However, this comes at the cost of slower inference speed due to the sequential nature of the interaction.
Figure 9: Coding collaboration with one-round discussion.
Discussion or debate frameworks (Figure 9) aim to improve response quality by enabling agents to access each other’s previous outputs (dudebate; liangdebate). In the first turn, the helper and main agents generate responses in the same manner as the naive concatenation approach. These initial responses are then shared with each other as references for the following turns, forming a discussion. Although this setup introduces more interaction, it does not guarantee improved response quality. With a limited number of rounds, the final output may not converge to a coherent solution. Even with additional rounds, convergence remains uncertain. This approach can even be less efficient than the sequential pipeline, particularly in distributed systems where communication latency is high or unreliable.
Figure 10: Coding collaboration in our method.
The interaction process between 2 agents trained with MAGRPO is illustrated in Figure 10. In the single-turn setting, we use the same prompts as in the naive concatenation baseline. In the multi-turn setting, after the helper and main agents generate their initial responses, these outputs are reviewed by an external agent. In this work, we employ Claude-Sonnet-4 as an external to provide edit suggestions. For each agent, the suggestions, as well as the prior information and their previous response, are incorporated into the prompt for the subsequent round.
Note that the baselines above can also be fine-tuned by MARL. However, the interactions among agents in these settings are not strictly cooperative, which may lead to instability during training. To address this, techniques such as role-based rewards (spiral), partial MAS training (trainaleader), and freezing selected agents (maft) can be employed to ensure stable optimization.
Appendix F Broader Impacts
@@
Prompt-based coordination is often brittle (debatefail), as agents may fail to follow instructions they were not explicitly trained to interpret. Our method builds on a solid theoretical foundation in cooperative MARL, explicitly optimizing agents for joint optimality. Our work also opens opportunities to enhance existing test-time multi-agent interaction methods by integrating MARL techniques (dudebate; duverifier; autogen), particularly in settings that involve task decomposition and iterative feedback integration.
@@
This work also explores a new perspective on accelerating LLM inference through cooperative MARL. While mainstream acceleration techniques (e.g., knowledge distillation, pruning, and quantization) improve efficiency at the cost of information loss (KDinformationloss; pruninginfoloss), our approach suggests decentralized coordination among specialized agents, thereby alleviating the burden of long-context memory and joint decision-making on a single model. Each agent can focus on a specific subtask, enabling more modular and robust reasoning.
Appendix G Limitations and Future Works
Nevertheless, this study is subject to several limitations:
  1. 1.
@@ Appendix G Limitations and Future Works
    First, we focus on homogeneous agents for simplicity, assuming they perform similar tasks despite being assigned different roles, e.g., both the auxiliary agent and main agent are generating Python functions. Future research could explore LLM collaboration among heterogeneous agents with diverse capabilities and functionalities.
  2. 2.
@@
    Due to computational constraints, we train LLMs with MAGRPO on limited datasets using relatively small-scale language models in a short horizon. When LLM-based coding agents are deployed in larger-scale projects involving multiple files and modules, more diverse and complex cooperation schemes are likely to emerge, which would further demonstrate the potential of decentralized coordination in MAS.
  3. 3.
@@
    The simplicity of our reward model inevitably leads to narrow reward signals and potential reward hacking. As suggested by many research studies and industrial practice (prm; multi-aspect; anthropic2023collective), designing more expressive and fine-grained reward models (e.g., multi-aspect rewards, process-supervised rewards) is essential for better aligning agent cooperation with human preferences.
Appendix H Compute Resources
We use NVIDIA H200 and H100 GPUs for LLM fine-tuning, and a standalone NVIDIA GeForce RTX 5090 workstation for inference. Here are the specifications of the resources used for experiments.
Training Devices
Type: GPU Cluster
CPU: Intel Xeon Platinum 8558
GPU: 1x NVIDIA H200
Inference Device
Type: Standalone Workstation
CPU: AMD Ryzen 9 9950X
GPU: 2x NVIDIA GeForce RTX 5090
The training process requires approximately 5 hours to complete 700 steps and 8 hours to complete 1000 steps. The inference takes approximately 0.5 hours for each 15-generation evaluation. Noted that training duration may vary considerably due to the node condition and the stochasticity of LLM outputs.
Appendix I Codes and Datasets
Codes
We developed an open-source library, CoMLRL, for training multiple LLMs to collaborate using MARL. CoMLRL provides configurable implementations of various MARL algorithms for LLM collaboration, including MAGRPO. The experiments of this paper serve as parts of CoMLRL’s environments and benchmarks.
  • OpenMLRL/CoMLRL:
    https://github.com/OpenMLRL/CoMLRL
  • OpenMLRL/LLM_Collab_Writing:
    https://github.com/OpenMLRL/LLM˙Collab˙Writing
  • OpenMLRL/LLM_Code_Generation:
    https://github.com/OpenMLRL/LLM˙Collab˙Code˙Generation
  • OpenMLRL/LLM_Code_Completion:
    https://github.com/OpenMLRL/LLM˙Collab˙Code˙Completion
We hope that this library and the associated experiment repositories will lay a foundation and provide convenient tools for future research and development in LLM collaborations.
Datasets
The datasets used in our experiments are all open-sourced and available online,
  • trl-lib/TLDR:
    https://huggingface.co/datasets/trl-lib/tldr
  • mattbierbaum/arXiv:
    https://github.com/mattbierbaum/arxiv-public-datasets
  • OpenAI/humanevalHumanEval:
    https://huggingface.co/datasets/openai/openai˙humaneval
  • OpenMLRL/CoopHumanEval:
    https://huggingface.co/datasets/OpenMLRL/CoopHumanEval